The following data for the Fufkin Company pertain to the produc­tion of 3,000 clay pigeons during July: Standard variable-overhead cost was $5.00 per pound of clay. The actual variable-overhead cost was $11,340. Standard variable-overhead cost allowed for units produced was $12,000. Variable-overhead efficiency variance was $340 unfavorable. _____ is the variable-overhead rate variance.
Answer #1

Variable overhead efficiency variance = (Standard hour-actual hour)Standard rate

-340 = (12000-5X)

5X = 12340

X(actual hour) = 2468 Hour

Variable overhead rate variance = (5*2468-11340)= 1000 F

So answer is $1000 F